the length of a rectangle is 5 meters less than twice its width.If the perimeter of the rectangle is 38 meters, then what is the value of the length(in meters)
Answer by checkley77(12569)

L=2W-5
2L+2W=38
2(2W-5)+2W=38
4W-10+2W=38
6W=38+10
6W=48
W=48/6
W=8 ANS. FOR THE WIDTH.
L=2*8-5
L=16-5
L=11 ANS. FOR THE LENGTH.
PROOF:
2*11+2*8=38
22+16=38
38=38